Solve the problem below and identify what gas law is used to solve it.

A 1.00 mol sample of gas at 25 °C and 1.0 atm has an initial volume of 22.4 L. Calculate the results of each change, assuming all the other conditions remain constant.
a. The pressure is changed to 85.7 mmHg. How many milliliters does the gas occupy?
b. The volume is reduced to 275 mL. What is the pressure in millimeters of mercury?à ƒ ¢ € ‹
1
Expert's answer
2021-10-19T10:13:01-0400

PV= nRT

0.00131579× 85.7= 0.113atm

0.113× V= 1× 8.314× 298

0.113V= 21925.42ml